OpenStudy (prettygirl_shynice):

may someone help with math i need help explaing like how i got answer i know the answer the answer but how i explaini know the answer its a but how i explain x to 50 and 16 to 25 32 34 41

zepdrix (zepdrix):

So you want to show what value of x makes `x to 50` equivalent to `16 to 25`. To get from 25 to 50, we double, yes? Multiply by 2. So that is our scale factor which relates the two ratios. So to get from 16 to x, we double :)

zepdrix (zepdrix):

That's a little bit of an explanation, hmm hopefully it makes sense to you..

OpenStudy (prettygirl_shynice):

i think so lol

OpenStudy (prettygirl_shynice):

omm may u explain a little eazier plz

OpenStudy (prettygirl_shynice):

srry

zepdrix (zepdrix):

This is the relationship the second numbers have: 50 = 2*25 So the first numbers have to follow that pattern as well: x = 2*16

OpenStudy (prettygirl_shynice):

2 time 16 is32

OpenStudy (prettygirl_shynice):

i follw the frist numbers

zepdrix (zepdrix):

so x is 32 :) yes `32 to 50` is equivalent to `16 to 25`

OpenStudy (prettygirl_shynice):

so frist i35 is equicwnt to 16 to 25

OpenStudy (prettygirl_shynice):

then i multiply 16 time 2 ?

OpenStudy (prettygirl_shynice):

@zepdrix am i doing it right

zepdrix (zepdrix):

If you prefer, you can think of these as fractions as opposed to ratios,\[\large\rm \frac{x}{50}=\frac{16}{25}\]And you can cross-multiply,\[\large\rm 25x=16\cdot50\]and then divide by 25,\[\large\rm x=\frac{16\cdot50}{25}\]

zepdrix (zepdrix):

Sorry, I'm having trouble understanding what you're saying :)

OpenStudy (prettygirl_shynice):

its okay me to

OpenStudy (prettygirl_shynice):

how u wright like that on the computer

zepdrix (zepdrix):

Below the text box, bottom left corner says "Equation". Press that button, it has a bunch of useful tools, let's you write out math in a fancy way.

OpenStudy (prettygirl_shynice):

no im talking about on word 2016

zepdrix (zepdrix):

On Microsoft Word? Hmm I'm not sure :)

OpenStudy (prettygirl_shynice):

oh lol

OpenStudy (prettygirl_shynice):

and yea

OpenStudy (prettygirl_shynice):

ah can u epxlain step by step like 1st 2nd 3rd

OpenStudy (prettygirl_shynice):

im srry but im still confused

OpenStudy (prettygirl_shynice):

i think got it